A 0.2 kg object is suspended from a spring with a spring constant of k=10 n/m. The object moves with simple harmonic motion and has an amplitude of 0.08 m . What is the potential energy of the system when the object displacement is 0.08 m ?
Answer #1

given


spring constant K = 10 N/m


amplitude = A = 0.08 m


mass = 0.2 kg

potetial of a body in SHM = PE = 0.5*K*x^2


x = displacement of the body from the mean position

here x = 0.08 m

PE = 0.5*10*0.08^2 = 0.032 J   <-------answer